OpenStudy (anonymous):

A small metal bar whose initial temperature was 20 C is dropped intoa large container of boiling water. How long will it take the bar to reach 90 C if it is known that it's temperature increases 2 C in 1 secnd. Howlong will it ake the bar to reach 98 C

OpenStudy (anonymous):

\[T_m=100\] \[\frac{dT}{dt}=k(T-100)\]

OpenStudy (anonymous):

\[2=k(90-100)\] 2=-10k -.2=k \[\frac{dT}{dt}=k(T-100)\] ln(T-100)=-.2t+c \[T=100+ce^{-.2t}\]

OpenStudy (anonymous):

\[20=100+ce^{-.2t}\] -80=c

OpenStudy (anonymous):

\[98=100-80e^{-.2t}\] \[\frac{-2}{-80}=e^{-.2t}\] \[ln(.025)=-.2t\] \[\frac{ln(.025)}{-.2}=t\]

OpenStudy (anonymous):

did icorrectly use the information to find k, since \[\frac{dT}{dt}=2\]

OpenStudy (anonymous):

the answer is like 86 seconds and i got like 16

jimthompson5910 (jim_thompson5910):

hmm your answer makes sense, does it give you any formulas you have to work with?

OpenStudy (anonymous):

only newtons law of cooling

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (anonymous):

i'm just confused allout by this because if it increases two degrees per seconds from 20 degrees it's a difference of 70 degrees

OpenStudy (anonymous):

common sense woul say that it would take half that to get it to 90 degrees

OpenStudy (anonymous):

which would be 35 seconds which is still no where near 82 seconds

OpenStudy (anonymous):

unless for some reason the temperature has tobe in farenheit

jimthompson5910 (jim_thompson5910):

hmm not too sure about this one either, but I did manage to find this http://faculty.frostburg.edu/math/rcforsythe/Math%20432/Solutions/Exam%202,%20Spring%202011%20--%20Solutions.pdf It's #2 in the pdf

OpenStudy (anonymous):

yeah that one is like most in the book... this one has the rate in it though =/

OpenStudy (anonymous):

it seems i have the rigt equation just the answer is wrong?

jimthompson5910 (jim_thompson5910):

I think it's the same problem, just worded differently

OpenStudy (anonymous):

ahh it is.. wow thanks lol

OpenStudy (anonymous):

that's a weird way to word a problem though. makes sense

jimthompson5910 (jim_thompson5910):

that's great it does